# Understanding World or Predicting Future? A Comprehensive Survey of World Models

### v2

@@ 3.2.3. Knowledge of the Human Society
Beyond the physical world, understanding human society is another crucial aspect of world models. David Premack and Guy Woodruff proposed the Theory of Mind (premack1978does, ), which was later developed to explain how individuals infer the mental states of others around them. Recent works have extensively explored how large language models develop and demonstrate this social world model (sap2022neural, ; strachan2024testing, ). Sap et al. (sap2022neural, ) conduct an investigation focusing on evaluating the performance of large language models across various Theory of Mind tasks to determine whether their human-like behaviors reflect genuine comprehension of social rules and implicit knowledge. Strachan et al. (strachan2024testing, ) conduct a comparative analysis between human and LLM performance on diverse Theory of Mind abilities, such as understanding false beliefs and recognizing irony. While their findings demonstrate the potential of GPT-4 in these tasks, they also identify its limitations, particularly in detecting faux pas.
To address these limitations, researchers propose innovative methods to enhance the abilities of large language models in Theory of Mind for complex real-world applications. Wu et al. (wu-etal-2024-coke, ) introduce COKE, which constructs a knowledge graph to help large language models explicitly using theory in mind through cognitive chains. Additionally, Alex et al. (wilf2023thinktwiceperspectivetakingimproves, ) develop SimToM, a two-stage prompting framework, to enhance the performance of large language models in theory of mind tasks.
@@ 4.1.1. Towards Video World Models
Sora as a World Simulator. Sora (sora2024, ) is a large-scale video generation model, which is designed to generate high-quality, temporally consistent video sequences, up to one minute long, based on various input modalities such as text, images, and videos. Sora leverages a combination of powerful neural network architectures, including encoder-decoder frameworks and transformers, to process multimodal inputs and generate visually coherent simulations. Sora’s core capabilities lie in its ability to generate videos that align with real-world physical principles, such as the reflection of light on surfaces or the melting of candles. These properties suggest that Sora has the potential to act as a world simulator, predicting future states of the world based on its understanding of the initial conditions and simulation parameters.
Sora’s Limitations. However, despite its impressive video generation abilities, Sora has several limitations in terms of understanding and simulating the external world. One key limitation concerns causal reasoning (zhu2024sora, ; cho2024sora, ), wherein the model is limited in simulating dynamic interactions within the environment. Thus, Sora can only passively generate video sequences based on an observed initial state, but cannot actively intervene or predict how changes in actions might alter the course of events. Another limitation is that it still fails to reproduce correct physical laws consistently (kang2024far, ). While Sora can generate visually realistic scenes, it struggles with accurately simulating real-world physics, such as the behavior of objects under different forces, fluid dynamics, or the accurate depiction of light and shadow interactions.
Other Video World Models. Sora has undoubtedly catalyzed a significant wave of research into video world models, inspiring a surge of advancements in this field. Following Sora’s success in generating high-quality video sequences, numerous subsequent models have been developed, each aiming to push the boundaries of what video world models can achieve. For example, some approaches have extended video lengths to enable long-form video simulation (yin2023nuwa, ; liu2024world, ; hu2023gaia, ). In addition to conventional language-guided video generation, more modalities are being integrated, such as images and actions (zhen20243d, ; xiang2024pandora, ). Researchers are also shifting their focus from basic video generation, which lacks user control, to interactive simulations that aim to replicate the decision space of the real world and facilitate decision-making (yang2023learning, ; yang2024video, ; wu2024ivideogpt, ; zhang2025physdreamer, ; jain2024peekaboo, ; xiang2024pandora, ). Several studies have worked to enhance the smoothness of action transitions, improve the accuracy of physical laws, and maintain temporal consistency (yang2024worldgpt, ; cai2023diffdreamer, ; ren2024consisti2v, ; yan2023temporally, ). Meanwhile, the concept of world models has evolved beyond imagination and is being applied in various scenario-specific simulations, including natural environments, games, and autonomous driving (liu2024world, ; wang2024worlddreamer, ; bruce2024genie, ; mendonca2023structured, ; hu2023gaia, ; wang2023drivedreamer, ; bogdoll2023muvo, ; zheng2025occworld, ; min2023uniworld, ). Table 2 summarizes the categorization of improvements in video world models across different aspects.
@@
PhysDreamer (zhang2025physdreamer, ) Simulates 3D object dynamics to generate responses to novel interactions. Diffusion
@@ 4.2.3. Dynamic Environments
Given the above developments, it is evident that world models as embodied environments have made significant advances in simulating and predicting how agents interact with dynamic, real-world scenarios. Current research predominantly focuses on developing indoor, static environments, with notable efforts expanding to large-scale outdoor environments and dynamic simulation environments. A promising direction is to construct dynamic environments, which can provide first-person, action-conditioned future world prediction, enabling agents to better adapt to unseen conditions. These methods are promising to offer flexible, scalable environments for training embodied agents, enhancing their generalization capabilities for real-world tasks.
5. Application
5.1. Autonomous Driving
@@
In recent years, with the rapid advancement of vision-based generative models (ho2020denoising, ; song2020score, ; videoworldsimulators2024, ) and multimodal large language models (liu2023llava, ; achiam2023gpt4, ), world models have attracted growing interest in the field of autonomous driving. The modern autonomous driving pipeline is typically divided into four key components: perception, prediction, planning, and control. Among these, the perception and prediction stages correspond to driving scene understanding—i.e., learning an implicit representation of the vehicle’s external environment. In parallel, recent surveys (guan2024world, ) highlight the emergence of end-to-end world simulators that learn to simulate realistic driving environments based on multimodal inputs—such as images, point clouds, trajectories, and language—and then generate future states to support downstream tasks like planning and decision-making. These two perspectives align well with our earlier categorization of world models, and in the following, we detail their applications and developments within the autonomous driving domain accordingly.
@@
5.1.1. Learning Implicit Representations
Autonomous vehicles typically utilize cameras, radar, and lidar to perceive the real world, gathering information through images, video data, and point cloud data. In the initial decision-making paradigm (chen2019modelfreedeepreinforcementlearning, ; Saxena_2020, ), models often take perceptual data as input and directly output motion planning results for the autonomous vehicle. Conversely, when humans operate vehicles, they typically observe and predict the current and future states of other traffic participants to determine their own driving strategies (survey_trajpre_2022, ). Thus, learning the implicit representation of the world through perceptual data and predicting the future states of the surrounding environment is a crucial step in enhancing the decision-making reliability of autonomous vehicles. We consider this process as it manifests in how autonomous vehicles learn a world model in latent space.
@@ 5.1.1. Learning Implicit Representations
5.1.2. World Simulators
As shown in Table 4, before the emergence of multimodal large models and vision-based generative models, traffic scenario simulations are often conducted in geometric spaces. The scene data on which these simulations rely is typically collected by the perception modules of autonomous vehicles or constructed manually. These simulations represent future states of the scenario in the form of geometric trajectories (SUMO2018, ; li2022metadrive, ; gulino2024waymax, ; zhang2023trafficbots, ), which require further modeling and rendering to produce outputs suitable for vehicle perception. The cascading of multiple modules often results in information loss and increases the complexity of simulations, making scenario control more challenging. Furthermore, realistic scene rendering typically requires substantial computational resources, which limits the efficiency of virtual traffic scenario generation.

### v4

@@ 3.2.3. Knowledge of the Human Society
Beyond the physical world, understanding human society is another crucial aspect of world models. David Premack and Guy Woodruff proposed the Theory of Mind (premack1978does, ), which was later developed to explain how individuals infer the mental states of others around them. Recent works have extensively explored how large language models develop and demonstrate this social world model (sap2022neural, ; strachan2024testing, ; kosinski2024evaluating, ). Sap et al. (sap2022neural, ) conduct an investigation focusing on evaluating the performance of large language models across various Theory of Mind tasks to determine whether their human-like behaviors reflect genuine comprehension of social rules and implicit knowledge. Strachan et al. (strachan2024testing, ) conduct a comparative analysis between human and LLM performance on diverse Theory of Mind abilities, such as understanding false beliefs and recognizing irony. While their findings demonstrate the potential of GPT-4 in these tasks, they also identify its limitations, particularly in detecting faux pas.
Beyond inferring individual mental states, researchers are also investigating how LLMs model the broader, underlying rules of human society. For instance, Mozikov et al. (mozikov2024eai, ) investigated how emotional factors influence ethical judgment and decision-making in LLMs, underscoring the need for robust mechanisms to ensure consistent ethical standards. Other works have explored the capacity of these models to navigate the complexities of a globalized world. Yin et al. (yin2024safeworld, ), for example, evaluated LLMs on their ability to generate responses that are not only helpful but also culturally sensitive and legally compliant across diverse global contexts. Similarly, Vayani (vayani2025all, ) conducted a large-scale evaluation of LLMs across a hundred culturally diverse languages, highlighting the importance of linguistic diversity in the development of truly global social models.
While these studies demonstrate the great potential of LLMs in modeling the social world, they also reveal significant limitations when these models must handle complex social situations. To address these shortcomings and enhance LLMs’ Theory of Mind abilities for complex, real-world applications, researchers have proposed several innovative methods. For instance, Wu et al. (wu-etal-2024-coke, ) introduced COKE, a framework that constructs a knowledge graph to help LLMs explicitly apply Theory of Mind through cognitive chains. Additionally, Alex et al. (wilf2023thinktwiceperspectivetakingimproves, ) developed SimToM, a two-stage prompting framework designed to improve the performance of large language models on Theory of Mind tasks.
@@ 4.1.1. Towards Video World Models
Sora (sora2024, ) is a large-scale video generation model designed for high-quality, temporally consistent video sequences up to one minute long, based on various multimodal inputs. It leverages neural network architectures to produce visually coherent simulations that often align with real-world physical principles, like light reflection or melting. These capabilities suggest Sora’s potential as a world simulator, predicting future states based on initial conditions and parameters. However, despite its impressive video generation, Sora has significant limitations in fully understanding and simulating the external world. A key limitation is its causal reasoning ability (zhu2024sora, ; cho2024sora, ), restricting it to passively generating sequences without actively predicting how actions might alter events. Furthermore, Sora struggles to consistently reproduce correct physical laws (kang2024far, ), failing to accurately simulate complex physics like object behavior under forces, fluid dynamics, or light interactions.
Following the success of Sora in generating high-quality videos, the past two years have witnessed the emergence of several large-scale video generative base models, such as OpenSora (zheng2024open, ), CogVideoX (yang2024cogvideox, ), and Wan (wan2025wan, ). These models, through the pre-training of more efficient VAEs and extensive pre-training on large-scale video datasets, exhibit strong visual generation capabilities and serve as foundational components for world models. Further advancing this field, Cosmos (agarwal2025cosmos, ) introduces a dedicated video generation base model for physical world simulation, achieving new breakthroughs in physical law adherence and understanding by pre-training on massive real-world physics videos and exploring both diffusion and auto-regressive architectures. Concurrently, Genie 2 (parkerholder2024genie2, ) and Genie 3 (genie3, ) focus on video generation in gaming scenarios, with Genie 2 specifically designing an auto-regressive diffusion architecture for interactive video generation that supports following external action instructions. Beyond these specific models, continuous progress is being made in key technical challenges including long-duration generation (yin2023nuwa, ; liu2024world, ; hu2023gaia, ; henschel2025streamingt2v, ), interactive generation (zhen20243d, ; xiang2024pandora, ; yang2023learning, ; yang2024video, ; wu2024ivideogpt, ; zhang2025physdreamer, ; jain2024peekaboo, ; xiang2024pandora, ; team2025aether, ; mao2025yume, ; bar2025navigation, ), and physical law adherence (yang2024worldgpt, ; cai2023diffdreamer, ; ren2024consisti2v, ; shang2025roboscape, ). Researchers are increasingly shifting focus from basic, user-uncontrolled video generation towards interactive simulations that replicate real-world decision spaces to facilitate decision-making. Moreover, the concept of world models has expanded beyond pure imagination, finding application in diverse scenario-specific simulations (liu2024world, ; wang2024worlddreamer, ; bruce2024genie, ; mendonca2023structured, ; hu2023gaia, ; wang2023drivedreamer, ; bogdoll2023muvo, ; min2023uniworld, ), encompassing natural environments, games, autonomous driving, and robotics.
@@
StreamingT2V (henschel2025streamingt2v, ) An autoregressive text-to-video model equipped with long/short-term memory blocks. Diffusion
@@
Aether (team2025aether, ) Utilizes camera trajectories as geometry-aware actions, enabling accurate action-conditioned prediction and visual planning. Diffusion
Yume (mao2025yume, ) Enhances streaming interactive world generation following continuous keyboard inputs. Diffusion
NWM (bar2025navigation, ) Achieves controllable video generation based on past observations and navigation actions. Diffusion
@@
WorldMem (xiao2025worldmem, ) Enhances long-term consistent world simulation with an integrated memory mechanism. Diffusion
@@ 4.2. World Model as Embodied Environment
Dynamic Roboscape (shang2025roboscape, ) Home 2025 Arbitrary V, T
Dynamic TesserAct (zhen2025tesseract, ) Home 2025 Arbitrary V, T
Dynamic Aether (team2025aether, ) Home, Street View 2025 Arbitrary V, T
Dynamic Deepverse (chen2025deepverse, ) Home, Simulation, Street View 2025 Arbitrary V, T
@@ 4.2.3. Dynamic Environments
Given the above developments, it is evident that world models as embodied environments have made significant advances in simulating the environmental transition of the real world. Current research predominantly focuses on developing indoor, static environments, with notable efforts expanding to large-scale outdoor and dynamic simulation environments. A promising direction is to construct dynamic environments, which can provide first-person, action-conditioned future world prediction, enabling agents to better adapt to unseen conditions. Simultaneously, recent advancements in building dynamic embodied worlds emphasize the integration of physical constraints. For instance, Aether (team2025aether, ) enhances geometric knowledge learning by employing camera trajectories as action-driven RGB-D video generation. TesserAct (zhen2025tesseract, ) further incorporates normal maps as physical constraints for video generation. Roboscape (shang2025roboscape, ) integrates depth maps and keypoint dynamics during video generation to learn and produce more realistic motions and spatial structures. Additionally, Deepverse (chen2025deepverse, ) proposes incorporating geometric predictions from previous timesteps into current predictions conditioned on actions. These methods collectively enhance the realism and physical adherence of the dynamic world, thereby creating more authentic and reliable simulated environments for embodied agents.
5. Application Domains
5.1. Game Intelligence
Gaming environments represent an ideal testbed for world model research, offering controlled yet complex domains that demand a sophisticated understanding of physics, causality, and interactive dynamics. Unlike real-world applications where ground truth is often ambiguous or inaccessible, games provide well-defined rule systems and clear action-consequence relationships that enable precise evaluation of world model capabilities.
More importantly, world model technologies are fundamentally transforming game development and player experiences in unprecedented ways. Traditional game development relies on manually coded rules, pre-designed assets, and scripted interactions that constrain creative possibilities and require extensive development resources. World models offer a paradigm shift toward generative game systems that can autonomously create new content, adapt dynamically to player behavior, and enable emergent gameplay experiences that were previously impossible to achieve through conventional programming approaches.
Recent developments demonstrate three critical capability dimensions in gaming applications:
Interactivity. The ability to respond appropriately to user inputs represents a fundamental requirement for gaming world models. GameNGen demonstrates this capability by creating a fully neural game engine that enables real-time interaction with complex environments, running at 20 frames per second while maintaining stable gameplay over extended sessions (valevski2024diffusion, ). Similarly, GameGen-X introduces a specialized module to incorporate game-related multi-modal control signals, unifying character interaction and scene content control for the first time in video generation (che2024gamegen, ). Matrix-Game advances this further by training an over 17 billion parameter model capable of precise control over both character actions and camera movements through fine-grained keyboard and mouse action annotations (zhang2025matrix, ).
Consistency. Maintaining coherent game states across temporal sequences poses significant challenges for generative models. Recent research tackles both numerical consistency (ensuring gameplay mechanics correctly reflect score changes and quantitative elements) and spatial consistency (preventing jarring scene transitions) in games (chen2025model, ). MineWorld addresses consistency through visual-action autoregressive transformers that learn rich representations of game states and action-state relationships simultaneously (guo2025mineworld, ). The model’s parallel decoding algorithm enables real-time generation while maintaining temporal coherence across extended gameplay sequences. WHAM (World and Human Action Model) further exemplifies this advancement by generating consistent and diverse gameplay sequences while persisting user modifications—capabilities identified as critical for supporting creative practices in game development (kanervisto2025world, ).
Generalization Across Diverse Environments. The ability to adapt across varied gaming scenarios and environments represents perhaps the most challenging dimension. GameFactory addresses this through scene-generalizable action control, leveraging open-domain generative priors from pre-trained video diffusion models to create entirely new games beyond fixed styles and scenes (yu2025gamefactory, ). Recent work further creates a ”generative infinite game” that transcends traditional finite, hard-coded systems (li2024unbounded, ). Using specialized distilled LLMs for dynamic game mechanic generation and dynamic regional image prompt adapters for consistent visual generation, the system enables open-ended mechanics that can emerge naturally from the underlying generative models. The exploration-driven approach in virtual environments demonstrates another path toward generalization, where exploration agents rely entirely on world model uncertainty to deliver diverse training data, adapting easily to new environments without environment-specific rewards (savov2025exploration, ).
5.2. Embodied Intelligence
Embodied intelligence focuses on creating agents that can perceive, understand, and interact effectively with the complex physical world. A central challenge in this domain is equipping robots with the ability to reason about their environment’s dynamics to support robust, real-time decision-making. World models have emerged as a transformative paradigm that directly addresses this need, empowering robots with the critical ability to perceive, predict, and act effectively. This progress is driven in part by advances in neural architectures (vaswani2017attention, ; ho2020denoising, ) and learning algorithms (schulman2017proximal, ; rafailov2024direct, ), which enable robots to build implicit representations that capture key aspects of the external world. Complementarily, prediction models (finn2016unsupervised, ; finn2017deep, ) offer the ability to forecast future environmental states, moving beyond static abstractions to support anticipatory and adaptive behavior. Together, these capabilities make it increasingly feasible for robots to learn directly from real-world interactions. In Table 4, we summarize the core learning tasks involved in constructing world models for robotics, categorized according to the three major perspectives outlined above (typical examples shown in Figure  S1).
5.2.1. Learning Implicit Representation
Task Model Year Input Backbone
ViT (dosovitskiy2020image, ) 2020 Image Transformer
RoboCraft (shi2024robocraft, ) 2024 Image GNN
Predictive Coding (gornet2024automated, ) 2024 Image ResNet
SpatialLM (spatiallm, ) 2025 3D point clouds MLLM
Task Representation BC-Z (jang2022bc, ) 2022 Text& Video LLM& ResNet
Text2Motion (lin2023text2motion, ) 2023 Text LLM
Gensim (wang2023gensim, ) 2023 Text LLM
VIPER (escontrela2024video, ) 2024 Video Transformer
GR-2 (cheang2024gr, ) 2024 Text & Video Transformer
IRASim (zhu2024irasim, ) 2024 Trajectory Diffusion
VPP (hu2024video, ) 2024 Text Diffusion
DreamGen (jang2025dreamgen, ) 2025 Text Diffusion
Roboscape (shang2025roboscape, ) 2025 Trajectory Transformer
EVAC (jiang2025enerverse, ) 2025 Trajectory Diffusion
Genie Envisioner (liao2025genie, ) 2025 Text Diffusion
Vidar (feng2025generalist, ) 2025 Text Diffusion
V-JEPA 2 (assran2025v, ) 2025 Trajectory Transformer
SWIM (mendonca2023structured, ) 2023 Video Transfer Learning
CoSTAR (agha2021nebula, ) 2021 Multimodal Belief Space
Evaluation OpenEQA (majumdar2024openeqa, ) 2024 Image& Text LLM
5.2.2. Predicting Future States of the Environment
World models are at the forefront of robotic research, primarily enabling advancements across three application areas: synthetic data generation, action guidance through imagined futures, and environment simulation for policy evaluation.
First, embodied world models can synthesize high-quality robot action videos to augment real-world collected data, thereby enhancing the training of downstream robotic policy models. For instance, DreamGen (jang2025dreamgen, ) proposes a four-stage pipeline to generate neural trajectories, a form of synthetic robot data derived from video world models. This approach significantly boosts robotic operation success and generalization in Vision-Language-Action (VLA) model training, particularly for contact-rich tasks. Roboscape (shang2025roboscape, ) integrates physical laws during video generation, leading to synthetic data with improved motion plausibility and spatial accuracy, which in turn yields substantial gains when incorporated into VLA training. EVAC (jiang2025enerverse, ) improves generalization by expanding training data with diverse failure trajectories, utilizing a multi-level action-conditioning mechanism and ray map encoding for dynamic multi-view image generation, effectively serving as both a data engine and an evaluator by augmenting human-collected trajectories.
Second, embodied world models guide robot action generation by leveraging imagined future observations. A key recent insight is the use of generative video models—particularly those leveraging diffusion (esser2023structure, ; chi2023diffusion, ; black2023zero, ; helearning, ) and transformer architectures (yu2023magvit, ; yan2021videogpt, )—to implicitly learn environmental dynamics directly from visual data. For instance, UniPi (du2024learning, ) explicitly frames action prediction as a video generation problem, conditioning a constrained diffusion model on the current state to visualize future scenarios. Similarly, VIPER (escontrela2024video, ) employs a pretrained autoregressive transformer to guide robotic actions, effectively leveraging rich representations learned from expert demonstration videos. Moreover, models like GR-2 (cheang2024gr, ) benefit from the vast scale of internet videos to establish robust priors, subsequently fine-tuning on specific robotics tasks to generate accurate image predictions and action trajectories. VPP (hu2024video, ) learns robot actions by fine-tuning a video generation model based on text instructions, subsequently deriving actions from an inverse dynamics model conditioned on visual representations. Similarly, Genie Envisioner (liao2025genie, ) leverages a pretrained embodied video generation foundation model, connected with a lightweight parallel flow-matching action model that translates language-conditioned visual latent features into fine-grained, low-latency motor commands. Vidar (feng2025generalist, ) proposes a two-stage framework for robotic action prediction, combining large-scale, diffusion-based video pre-training with a novel masked inverse dynamics model. Distinctly, V-JEPA 2 (assran2025v, ) models world state transitions in a latent space and performs action planning through Model Predictive Control (MPC), which involves extensive sampling of possible action trajectories and selecting the optimal one based on energy optimization.
Third, embodied world models can function as environment simulators for policy evaluation. IRASim (zhu2024irasim, ) and Roboscape (shang2025roboscape, ) both leverage world models for trajectory-to-video generation tasks, starting from an initial given frame. Their demonstrated high correlation between policy model evaluation in the world model and in real environments indicates that the learned world models accurately capture world transition dynamics. GE-Sim (liao2025genie, ) also exemplifies this by designing a closed-loop interaction between a policy and a world model, enabling scalable and flexible simulation without requiring manual environment modeling.
Collectively, these methods demonstrate the profound promise of generative, vision-centric world modeling as a foundation for anticipatory robotic control and simulation. This significantly enhances robots’ capabilities to reason about future states and improve long-term task performance.
5.2.3. From Simulation to Real World
5.3. Urban Intelligence
5.3.1. Autonomous Driving
@@
In recent years, with the rapid advancement of vision-based generative models (ho2020denoising, ; song2020score, ; videoworldsimulators2024, ) and multimodal large language models (liu2023llava; achiam2023gpt4, ), world models have attracted growing interest in the field of autonomous driving. The modern autonomous driving pipeline is typically divided into four key components: perception, prediction, planning, and control. Among these, the perception and prediction stages correspond to driving scene understanding—i.e., learning an implicit representation of the vehicle’s external environment. In parallel, recent surveys (guan2024world, ) highlight the emergence of end-to-end world simulators that learn to simulate realistic driving environments based on multimodal inputs—such as images, point clouds, trajectories, and language—and then generate future states to support downstream tasks like planning and decision-making. These two perspectives align well with our earlier categorization of world models, and in the following, we detail their applications and developments within the autonomous driving domain accordingly.
@@
Learning Implicit Representations. Autonomous vehicles typically utilize cameras, radar, and lidar to perceive the real world, gathering information through images, video data, and point cloud data. In the initial decision-making paradigm (chen2019modelfreedeepreinforcementlearning, ; Saxena_2020, ), models often take perceptual data as input and directly output motion planning results for the autonomous vehicle. Conversely, when humans operate vehicles, they typically observe and predict the current and future states of other traffic participants to determine their own driving strategies (survey_trajpre_2022, ). Thus, learning the implicit representation of the world through perceptual data and predicting the future states of the surrounding environment is a crucial step in enhancing the decision-making reliability of autonomous vehicles. We consider this process as it manifests in how autonomous vehicles learn a world model in latent space.
@@
World Simulators. As shown in Table 5, before the emergence of multimodal large models and vision-based generative models, traffic scenario simulations are often conducted in geometric spaces. The scene data on which these simulations rely is typically collected by the perception modules of autonomous vehicles or constructed manually. These simulations represent future states of the scenario in the form of geometric trajectories (SUMO2018, ; li2022metadrive, ; gulino2024waymax, ; zhang2023trafficbots, ), which require further modeling and rendering to produce outputs suitable for vehicle perception. The cascading of multiple modules often results in information loss and increases the complexity of simulations, making scenario control more challenging. Furthermore, realistic scene rendering typically requires substantial computational resources, which limits the efficiency of virtual traffic scenario generation.